# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v2

@@ 4 First Application with DESI Y1 BGS
In this section, we present the first application of the improved PAC method using the DESI Y1 BGS and DECaLS samples. We derive $\bar{n}_{2}w_{{\rm{p}}}$ for blue and red ${\rm{pop}}_{2}$ galaxies down to stellar masses of $10^{5.3}{\rm M}_{\sun}$ and $10^{6.3}{\rm M}_{\sun}$, respectively. Following the approach of Paper III, we obtain measurements of the GSMFs for blue and red galaxies at these mass thresholds by combining $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ from the BGS dataset. We then combine these to obtain the total GSMF.
@@ 4.1 Methodology for measuring GSMF
Paper III proposed a method to directly derive the galaxy GSMF by combining $\bar{n}_{2}w_{{\rm{p}}}$ from the PAC method with $w_{{\rm{p}}}$ from spectroscopic surveys, which may not be complete in stellar mass. In this approach, to match ${\rm{pop}}_{2}$ from the photometric sample, the corresponding ${\rm{pop}}_{2}^{\rm{s}}$ is selected from the spectroscopic sample within the same stellar mass bin and cross-correlated with ${\rm{pop}}_{1}$ to obtain $w_{{\rm{p}}}$. The GSMF is then estimated by comparing $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ within each stellar mass bin. This method is straightforward and relies on the key assumption that galaxy bias is primarily determined by stellar mass, such that ${\rm{pop}}_{2}^{\rm{s}}$ and ${\rm{pop}}_{2}$ have similar biases, even if ${\rm{pop}}_{2}^{\rm{s}}$ is highly incomplete. In this paper, we use a similar approach with several improvements and extensions.
Although $w_{{\rm{p}}}$ is primarily determined by stellar mass, residual scatter has been observed in numerous studies, which has been shown to depend on secondary properties such as colour and morphology (Li et al., 2006; Wang et al., 2016a; Xu & Jing, 2022; Gao et al., 2022). For example, massive red central galaxies are found in more massive halos compared to massive blue ones, resulting in stronger clustering (Wang et al., 2016a; Xu & Jing, 2022). Moreover, red galaxies have a higher satellite fraction than blue galaxies (Gao et al., 2022). These red satellites reside in more massive halos, which enhances the clustering of red galaxies at small scales (1-halo term) and slightly increases it at large scales (2-halo term). Consequently, at a fixed stellar mass, red galaxies exhibit higher clustering than blue galaxies, with the difference becoming more pronounced on smaller scales (Li et al., 2006). Therefore, the method used in Paper III may overestimate the GSMF, as red galaxies, with their higher mass-to-light ratios, are often among the first to be excluded when the sample becomes incomplete. This bias skews the sample toward a higher fraction of blue galaxies, leading to lower observed $w_{{\rm{p}}}$.
@@
Additionally, we also divide ${\rm{pop}}_{1}$ into different stellar mass bins following Paper III. For a given ${\rm{pop}}_{2}$ sample, the value of $\bar{n}_{2}$ from each ${\rm{pop}}_{1}$ sample should be consistent, as the method is independent of the properties of ${\rm{pop}}_{1}$. Therefore, comparing results from ${\rm{pop}}_{1}$ across different stellar mass bins serves as a robust test of our method. Additionally, refining ${\rm{pop}}_{1}$ into smaller stellar mass bins makes the $\bar{n}_{2}w_{{\rm{p}}}$ and $w_{{\rm{p}}}$ measurements easier to interpret.
@@ 4.2 Measurements of $\bar{n}_{2}w_{{\rm{p}}}$
We measure $\bar{n}_{2}w_{{\rm{p}}}$ for ${\rm{pop}_{2}}$ from DECaLS over a wide stellar mass range of $[10^{6.3}{\rm M}_{\sun},10^{11.9}{\rm M}_{\sun}]$ for red galaxies and $[10^{5.3}{\rm M}_{\sun},10^{11.9}{\rm M}_{\sun}]$ for blue galaxies, using a bin width of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.2$, cross-correlated with ${\rm{pop}_{1}}$ from DESI BGS in 11 stellar mass bins within $[10^{6.5}{\rm M}_{\sun},10^{12.0}{\rm M}_{\sun}]$ and with a bin width of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.5$. The calculation is done across the redshift ranges where ${\rm{pop}_{2}}$ is complete, according to the results in Figure 6. The completeness of stellar mass for ${\rm{pop}_{1}}$ is not considered, as the results are independent of the properties of ${\rm{pop}_{1}}$, as mentioned in Section 4.1. Our analysis covers spatial scales ranging from $0.01\,h^{-1}\rm{Mpc}$ to $25\,h^{-1}\rm{Mpc}$, showcasing an additional advantage of the PAC method as it remains unaffected by fibre incompleteness, which is a main challenge for small scale measurements in spectroscopic surveys.
@@ 4.3 Measurements of $w_{{\rm{p}}}$
The cross-correlation functions $\xi(r_{\rm{p}},\pi)$ are estimated using the Landy-Szalay estimator (Landy & Szalay, 1993) within the corresponding $r_{\rm{p}}$ bins and 40 $\pi$ bins, ranging from 00 to $40\,h^{-1}\rm{Mpc}$ in equal linear intervals, computed with Corrfunc (Sinha & Garrison, 2020). The projected correlation functions $w_{{\rm{p}}}$ are then calculated as:
[…]
@@
For blue galaxies, we present the measurements of $w_{{\rm{p}}}$ for ${\rm{pop}}_{2}^{\rm{s}}$ across stellar mass bins ranging from $10^{6.8}{\rm M}_{\sun}$ to $10^{10.0}{\rm M}_{\sun}$ in Figure 8. Each panel shows the results for one of the 11 ${\rm{pop}_{1}}$ stellar mass bins. To illustrate the scale at which $w_{{\rm{p}}}$ may be impacted by fibre incompleteness, we include a dotted line at $0.05^{\circ}$ for $z=0.2$. As shown in Figure 1, low-mass galaxies may be found at lower redshifts, allowing for $w_{{\rm{p}}}$ measurements at smaller $r_{\rm{p}}$. The angle $0.05^{\circ}$ is selected as the patrol diameter of the fibre positioners extends up to $\sim 0.0495^{\circ}$. This choice is validated by tests in Smith et al. (2019). We find that blue galaxies have nearly identical $w_{{\rm{p}}}$ values at $M_{*}<10^{10.0}{\rm M}_{\sun}$. This can be explained by the fact that blue galaxies are predominantly central galaxies (Gao et al., 2022, 2023). Their correlation functions are largely driven by halo-halo correlation, with subhalo contributions being minimal. It is well-established that halo bias remains nearly constant for low-mass halos ($M_{\rm{h}}<10^{12.0}h^{-1}{\rm M}_{\sun}$) (Jing, 1998; Tinker et al., 2010), which could explain why low-mass blue galaxies show nearly the same correlation functions. To extend the GSMF for blue galaxies to $M_{*}<10^{6.3}{\rm M}_{\sun}$, for these mass bins, we measure the $w_{{\rm{p}}}$ from all the blue ${\rm{pop}}_{2}^{\rm{s}}$ with $M_{*}<10^{9.0}{\rm M}_{\sun}$, and assume that it is the $w_{{\rm{p}}}$ of ${\rm{pop}}_{2}$. To validate this approximation, we also apply it to ${\rm{pop}}_{2}$ with $10^{6.3}{\rm M}_{\sun}<M_{*}<10^{9.1}{\rm M}_{\sun}$ and compare the resulting GSMF with the fiducial results.
@@
For red galaxies, we present their $w_{{\rm{p}}}$ measurements across several stellar mass bins ranging from $10^{8.4}{\rm M}_{\sun}$ to $10^{10.8}{\rm M}_{\sun}$ in Figure 9. In contrast to blue galaxies, we observe a clear dependence of $w_{{\rm{p}}}$ on stellar mass, with a steeper slope at lower masses. This could be attributed to the fact that a significant fraction of low-mass red galaxies are satellites, and their fraction, host halo mass, and radial distribution may vary with stellar mass. As shown in Figure 9, lower mass red galaxies appear to have a higher satellite fraction, which contributes to the steeper $w_{{\rm{p}}}$ slope. Despite this stellar mass dependence, we find that $w_{{\rm{p}}}$ values for different stellar mass bins are similar at $r_{\rm{p}}\sim 10\,h^{-1}\rm{Mpc}$. This consistency is expected. At $8\,h^{-1}\mathrm{Mpc}$, $w_{{\rm{p}}}$ is dominated by the 2-halo term, meaning its amplitude is primarily determined by the mean host halo bias of the red galaxies. Halo bias changes slightly for $M_{\rm{h}}<10^{12}h^{-1}{\rm M}_{\sun}$ (0.6-0.8) and even up to $M_{\rm{h}}<10^{13}h^{-1}{\rm M}_{\sun}$ ($\sim 1.0$; Jing, 1998; Tinker et al., 2010). As long as most low-mass galaxies reside in halos with $M_{\rm{h}}<10^{13}h^{-1}{\rm M}_{\sun}$, $w_{{\rm{p}}}$ at $8\,h^{-1}\mathrm{Mpc}$ should remain relatively constant across different stellar masses. This holds true for two reasons: (1) If they are centrals, they must be hosted by low-mass halos. (2) For satellites, the number density of halos with $M_{\rm{h}}>10^{13}h^{-1}{\rm M}_{\sun}$ decreases exponentially, so most satellites are expected to reside in relatively lower-mass host halos. To estimate $w_{{\rm{p}}}$ for red galaxies with $M_{*}<10^{8.1}{\rm M}_{\sun}$, we assume that their $w_{{\rm{p}}}$ at $8\,h^{-1}\rm{Mpc}$ is consistent with the $w_{{\rm{p}}}$ of all red ${\rm{pop}}_{2}^{\rm{s}}$ galaxies with $M_{*}<10^{10.8}{\rm M}_{\sun}$. Additionally, we assume that the correlation function follows a power-law form, $\xi=(r/r_{0})^{-\gamma}$, which implies that $w_{{\rm{p}}}$ also follows a power law:
@@
Additionally, in a companion paper, we plan to derive the GSMF for the entire sample by modelling their $\bar{n}_{2}w_{{\rm{p}}}$ using the subhalo abundance matching method within numerical simulations, as developed in Paper IV. This approach will further validate our results, as it does not rely on assumptions about $w_{{\rm{p}}}$ and is solely based on the $\bar{n}_{2}w_{{\rm{p}}}$ measurements.
@@ 4.4 Obtaining $\bar{n}_{2}$
With these measurements, we can calculate $\bar{n}_{2}$ by simply dividing $\bar{n}_{2}w_{{\rm{p}}}$ by $w_{{\rm{p}}}$. However, similar to Paper III, instead of employing error propagation, we estimate $\bar{n}_{2}$ in each sub-region and then combine them to get the final results:
@@
where $\chi^{2}_{m}$ is the $\chi^{2}$ for the $m$th stellar mass bin of pop1. We use the Markov chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to perform maximum likelihood analyses of $\{\bar{n}_{2}\}$ for each stellar mass bin of pop2.
@@ 4.5 Lower limits for smaller stellar masses
To achieve this, we first apply an $r$-band magnitude cut of 23.0 to the DECaLS data, a limit reached by over $90\%$ of the regions in DECaLS (see Paper III, ). This ensures a uniform sample for calculating $\bar{n}_{2}w_{{\rm{p}}}$ in the incomplete stellar mass bins. We then compute $\bar{n}_{2}w_{{\rm{p}}}$ for blue and red galaxies down to $10^{3.7}{\rm M}_{\sun}$ and $10^{5.5}{\rm M}_{\sun}$, respectively, by cross-correlating DECaLS with the DESI BGS sample within the redshift range $z<0.02$. Our analysis shows that we successfully obtain $\bar{n}_{2}w_{{\rm{p}}}$ measurements for these stellar mass bins. To determine the lower limits of $\bar{n}_{2}$ for these bins, we apply fixed $w_{{\rm{p}}}$ and $w_{{\rm{p}}}$ fitting methods for blue and red galaxies, respectively. The fits for blue and red galaxies are shown in Figures 26 and 27.
@@
Figure 15: Comparison to previous GSMFs measured in the nearby universe. The GSMF of the full sample (black dots) is compared to results from Paper III, obtained using the old PAC method applied to SDSS data (blue diamonds) at $z<0.2$, GAMA DR4 results (orange stars; Driver et al., 2022) at $z<0.1$, and the DESI BGS Bright sample analyzed with the $V_{\rm{max}}$ method (green dots) at $z<0.2$ (S. Moore 2025, in preparation). Systematic errors in stellar mass for the GAMA results have been corrected using the relationship shown in Figure 16. Additionally, DESI BGS results after correcting for the Local Void bias are shown (purple dots). This correction is based on Chen et al. (2019), leveraging the ELUCID simulation (Wang et al., 2014b, 2016b), which uses the reconstructed initial density field of the nearby universe. The GSMF from (Driver et al., 2022) for Local Group galaxies (brown triages; McConnachie, 2012) is also reproduced using the same stellar mass calibration as applied to GAMA.
@@
Figure 16: Left: Comparing the stellar mass $M_{*}^{\rm{GAMA}}$ obtained from GAMA DR4 (Driver et al., 2022) with those derived using our CIGALE SED template (described in Section 2.1) based on DECaLS $grz$ bands photometry ($M_{*}^{\rm{CIGALE}}$). The relationship between the two can be effectively characterized by a power law: $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}{\rm M}_{\sun})=1.072\log_{10}(M^{\rm{GAMA}}_{*}/10^{10}{\rm M}_{\sun})-0.237$ (red line). Right: Comparing the stellar mass $M_{*}^{\rm{GAMA}}$ obtained from COSMOS2020 (Weaver et al., 2023) with those derived using our CIGALE SED template based on HSC $grz$ bands photometry at $0.80.2<z_{\rm p}<$ 0.8. Black dots with errorbars represent the mean relation with the 1 $\sigma$ scatter. The relationship between the two can be effectively characterized by a power law: $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}{\rm M}_{\sun})=0.948\log_{10}(M^{\rm{COSMOS}}_{*}/10^{10}{\rm M}_{\sun})-0.012$.
@@
Figure 17: Top: Comparison with galaxy stellar mass functions (GSMFs) from the COSMOS field at higher redshifts. The GSMF of the full sample (black dots) is compared with results from COSMOS-Web (Shuntov et al., 2025) at $0.50.2<z_{\rm p}<$ 0.5 (blue dots) and $0.80.5<z_{\rm p}<$ 0.8 (pink dots). Stellar masses from COSMOS-Web have been calibrated to match our scale using the right panel of Figure 16. GSMFs from G10-COSMOS (Wright et al., 2017) at $z_{\rm p}<1$ are also shown (green triangles) with stellar mass calibrated with the relation for GAMA. Bottom: Stellar mass density within each stellar mass interval.

### v3

@@ 4 First Application with DESI Y1 BGS
In this section, we present the first application of the improved PAC method using the DESI Y1 BGS and DECaLS samples. We derive $\bar{n}_{2}w_{{\rm{p}}}$ for blue and red ${\rm{pop}}_{2}$ galaxies down to stellar masses of $10^{5.3}{\rm M}_{\sun}$ and $10^{6.3}{\rm M}_{\sun}$, respectively. Following the approach of 2022ApJ...939..104X, we obtain measurements of the GSMFs for blue and red galaxies at these mass thresholds by combining $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ from the BGS dataset. We then combine these to obtain the total GSMF.
@@ 4.1 Methodology for measuring GSMF
2022ApJ...939..104X proposed a method to directly derive the galaxy GSMF by combining $\bar{n}_{2}w_{{\rm{p}}}$ from the PAC method with $w_{{\rm{p}}}$ from spectroscopic surveys, which may not be complete in stellar mass. In this approach, to match ${\rm{pop}}_{2}$ from the photometric sample, the corresponding ${\rm{pop}}_{2}^{\rm{s}}$ is selected from the spectroscopic sample within the same stellar mass bin and cross-correlated with ${\rm{pop}}_{1}$ to obtain $w_{{\rm{p}}}$. The GSMF is then estimated by comparing $\bar{n}_{2}w_{{\rm{p}}}$ with $w_{{\rm{p}}}$ within each stellar mass bin. This method is straightforward and relies on the key assumption that galaxy bias is primarily determined by stellar mass, such that ${\rm{pop}}_{2}^{\rm{s}}$ and ${\rm{pop}}_{2}$ have similar biases, even if ${\rm{pop}}_{2}^{\rm{s}}$ is highly incomplete. In this paper, we use a similar approach with several improvements and extensions.
Although $w_{{\rm{p}}}$ is primarily determined by stellar mass, residual scatter has been observed in numerous studies, which has been shown to depend on secondary properties such as colour and morphology (2006MNRAS.368...21L; 2016ApJ...819...58W; 2022ApJ...926..130X; 2022ApJ...928...10G). For example, massive red central galaxies are found in more massive halos compared to massive blue ones, resulting in stronger clustering (2016ApJ...819...58W; 2022ApJ...926..130X). Moreover, red galaxies have a higher satellite fraction than blue galaxies (2022ApJ...928...10G). These red satellites reside in more massive halos, which enhances the clustering of red galaxies at small scales (1-halo term) and slightly increases it at large scales (2-halo term). Consequently, at a fixed stellar mass, red galaxies exhibit higher clustering than blue galaxies, with the difference becoming more pronounced on smaller scales (2006MNRAS.368...21L). Therefore, the method used in 2022ApJ...939..104X may overestimate the GSMF, as red galaxies, with their higher mass-to-light ratios, are often among the first to be excluded when the sample becomes incomplete. This bias skews the sample toward a higher fraction of blue galaxies, leading to lower observed $w_{{\rm{p}}}$.
@@
Additionally, we also divide ${\rm{pop}}_{1}$ into different stellar mass bins following 2022ApJ...939..104X. For a given ${\rm{pop}}_{2}$ sample, the value of $\bar{n}_{2}$ from each ${\rm{pop}}_{1}$ sample should be consistent, as the method is independent of the properties of ${\rm{pop}}_{1}$. Therefore, comparing results from ${\rm{pop}}_{1}$ across different stellar mass bins serves as a robust test of our method. Additionally, refining ${\rm{pop}}_{1}$ into smaller stellar mass bins makes the $\bar{n}_{2}w_{{\rm{p}}}$ and $w_{{\rm{p}}}$ measurements easier to interpret.
@@ 4.2 Measurements of $\bar{n}_{2}w_{{\rm{p}}}$
We measure $\bar{n}_{2}w_{{\rm{p}}}$ for $\rm{pop}_{2}$} from DECaLS over a wide stellar mass range of $10^{6.3}{\rm M}_{\sun},10^{11.9}{\rm M}_{\sun}]$ for red galaxies and $10^{5.3}{\rm M}_{\sun},10^{11.9}{\rm M}_{\sun}]$ for blue galaxies, using a bin width of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.2$, cross-correlated with $\rm{pop}_{1}$} from DESI BGS in 11 stellar mass bins within $10^{6.5}{\rm M}_{\sun},10^{12.0}{\rm M}_{\sun}]$ and with a bin width of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.5$. The calculation is done across the redshift ranges where $\rm{pop}_{2}$} is complete, according to the results in Figure 6. The completeness of stellar mass for $\rm{pop}_{1}$} is not considered, as the results are independent of the properties of $\rm{pop}_{1}$}, as mentioned in Section 4.1. Our analysis covers spatial scales ranging from $0.01\,h^{-1}\rm{Mpc}$ to $25\,h^{-1}\rm{Mpc}$, showcasing an additional advantage of the PAC method as it remains unaffected by fibre incompleteness, which is a main challenge for small scale measurements in spectroscopic surveys.
@@ 4.3 Measurements of $w_{{\rm{p}}}$
The cross-correlation functions $\xi(r_{\rm{p}},\pi)$ are estimated using the Landy-Szalay estimator (1993ApJ...412...64L) within the corresponding $r_{\rm{p}}$ bins and 40 $\pi$ bins, ranging from 0 to $40\,h^{-1}\rm{Mpc}$ in equal linear intervals, computed with Corrfunc (2020MNRAS.491.3022S). The projected correlation functions $w_{{\rm{p}}}$ are then calculated as:
$w_{{\rm{p}}}(r_{\rm{p}})=2\int_{0}^{\pi_{\rm{max}}}\xi(r_{\rm{p}},\pi)d\pi\,$, (22)
@@
For blue galaxies, we present the measurements of $w_{{\rm{p}}}$ for ${\rm{pop}}_{2}^{\rm{s}}$ across stellar mass bins ranging from $10^{6.8}{\rm M}_{\sun}$ to $10^{10.0}{\rm M}_{\sun}$ in Figure 8. Each panel shows the results for one of the 11 $\rm{pop}_{1}$} stellar mass bins. To illustrate the scale at which $w_{{\rm{p}}}$ may be impacted by fibre incompleteness, we include a dotted line at $0.05^{\circ}$ for $z=0.2$. As shown in Figure 1, low-mass galaxies may be found at lower redshifts, allowing for $w_{{\rm{p}}}$ measurements at smaller $r_{\rm{p}}$. The angle $0.05^{\circ}$ is selected as the patrol diameter of the fibre positioners extends up to $\sim 0.0495^{\circ}$. This choice is validated by tests in 2019MNRAS.484.1285S. We find that blue galaxies have nearly identical $w_{{\rm{p}}}$ values at $M_{*}<10^{10.0}{\rm M}_{\sun}$. This can be explained by the fact that blue galaxies are predominantly central galaxies (2022ApJ...928...10G; 2023ApJ...954..207G). Their correlation functions are largely driven by halo-halo correlation, with subhalo contributions being minimal. It is well-established that halo bias remains nearly constant for low-mass halos ($M_{\rm{h}}<10^{12.0}h^{-1}{\rm M}_{\sun}$) (1998ApJ...503L...9J; 2010ApJ...724..878T), which could explain why low-mass blue galaxies show nearly the same correlation functions. To extend the GSMF for blue galaxies to $M_{*}<10^{6.3}{\rm M}_{\sun}$, for these mass bins, we measure the $w_{{\rm{p}}}$ from all the blue ${\rm{pop}}_{2}^{\rm{s}}$ with $M_{*}<10^{9.0}{\rm M}_{\sun}$, and assume that it is the $w_{{\rm{p}}}$ of ${\rm{pop}}_{2}$. To validate this approximation, we also apply it to ${\rm{pop}}_{2}$ with $10^{6.3}{\rm M}_{\sun}<M_{*}<10^{9.1}{\rm M}_{\sun}$ and compare the resulting GSMF with the fiducial results.
@@
For red galaxies, we present their $w_{{\rm{p}}}$ measurements across several stellar mass bins ranging from $10^{8.4}{\rm M}_{\sun}$ to $10^{10.8}{\rm M}_{\sun}$ in Figure 9. In contrast to blue galaxies, we observe a clear dependence of $w_{{\rm{p}}}$ on stellar mass, with a steeper slope at lower masses. This could be attributed to the fact that a significant fraction of low-mass red galaxies are satellites, and their fraction, host halo mass, and radial distribution may vary with stellar mass. As shown in Figure 9, lower mass red galaxies appear to have a higher satellite fraction, which contributes to the steeper $w_{{\rm{p}}}$ slope. Despite this stellar mass dependence, we find that $w_{{\rm{p}}}$ values for different stellar mass bins are similar at $r_{\rm{p}}\sim 10\,h^{-1}\rm{Mpc}$. This consistency is expected. At $8\,h^{-1}\mathrm{Mpc}$, $w_{{\rm{p}}}$ is dominated by the 2-halo term, meaning its amplitude is primarily determined by the mean host halo bias of the red galaxies. Halo bias changes slightly for $M_{\rm{h}}<10^{12}h^{-1}{\rm M}_{\sun}$ (0.6-0.8) and even up to $M_{\rm{h}}<10^{13}h^{-1}{\rm M}_{\sun}$ ($\sim 1.0$; 1998ApJ...503L...9J; 2010ApJ...724..878T). As long as most low-mass galaxies reside in halos with $M_{\rm{h}}<10^{13}h^{-1}{\rm M}_{\sun}$, $w_{{\rm{p}}}$ at $8\,h^{-1}\mathrm{Mpc}$ should remain relatively constant across different stellar masses. This holds true for two reasons: (1) If they are centrals, they must be hosted by low-mass halos. (2) For satellites, the number density of halos with $M_{\rm{h}}>10^{13}h^{-1}{\rm M}_{\sun}$ decreases exponentially, so most satellites are expected to reside in relatively lower-mass host halos. To estimate $w_{{\rm{p}}}$ for red galaxies with $M_{*}<10^{8.1}{\rm M}_{\sun}$, we assume that their $w_{{\rm{p}}}$ at $8\,h^{-1}\rm{Mpc}$ is consistent with the $w_{{\rm{p}}}$ of all red ${\rm{pop}}_{2}^{\rm{s}}$ galaxies with $M_{*}<10^{10.8}{\rm M}_{\sun}$. Additionally, we assume that the correlation function follows a power-law form, $\xi=(r/r_{0})^{-\gamma}$, which implies that $w_{{\rm{p}}}$ also follows a power law:
@@
Additionally, in a companion paper, we plan to derive the GSMF for the entire sample by modelling their $\bar{n}_{2}w_{{\rm{p}}}$ using the subhalo abundance matching method within numerical simulations, as developed in 2023ApJ...944..200X. This approach will further validate our results, as it does not rely on assumptions about $w_{{\rm{p}}}$ and is solely based on the $\bar{n}_{2}w_{{\rm{p}}}$ measurements.
@@ 4.4 Obtaining $\bar{n}_{2}$
With these measurements, we can calculate $\bar{n}_{2}$ by simply dividing $\bar{n}_{2}w_{{\rm{p}}}$ by $w_{{\rm{p}}}$. However, similar to 2022ApJ...939..104X, instead of employing error propagation, we estimate $\bar{n}_{2}$ in each sub-region and then combine them to get the final results:
@@
where $\chi^{2}_{m}$ is the $\chi^{2}$ for the $m$th stellar mass bin of pop1. We use the Markov chain Monte Carlo sampler emcee (2013PASP..125..306F) to perform maximum likelihood analyses of $\{\bar{n}_{2}\}$ for each stellar mass bin of pop2.
@@ 4.5 Lower limits for smaller stellar masses
To achieve this, we first apply an $r$-band magnitude cut of 23.0 to the DECaLS data, a limit reached by over $90\%$ of the regions in DECaLS (see 2022ApJ...939..104X, ). This ensures a uniform sample for calculating $\bar{n}_{2}w_{{\rm{p}}}$ in the incomplete stellar mass bins. We then compute $\bar{n}_{2}w_{{\rm{p}}}$ for blue and red galaxies down to $10^{3.7}{\rm M}_{\sun}$ and $10^{5.5}{\rm M}_{\sun}$, respectively, by cross-correlating DECaLS with the DESI BGS sample within the redshift range $z<0.02$. Our analysis shows that we successfully obtain $\bar{n}_{2}w_{{\rm{p}}}$ measurements for these stellar mass bins. To determine the lower limits of $\bar{n}_{2}$ for these bins, we apply fixed $w_{{\rm{p}}}$ and $w_{{\rm{p}}}$ fitting methods for blue and red galaxies, respectively. The fits for blue and red galaxies are shown in Figures 26 and 27.
@@
Figure 15: Comparison to previous GSMFs measured in the nearby universe. The GSMF of the full sample (black dots) is compared to results from 2022ApJ...939..104X, obtained using the old PAC method applied to SDSS data (blue diamonds) at $z<0.2$, GAMA DR4 results (orange stars; 2022MNRAS.513..439D) at $z<0.1$, and the DESI BGS Bright sample analyzed with the $V_{\rm{max}}$ method (green dots) at $z<0.2$ (S. Moore 2025, in preparation). Systematic errors in stellar mass for the GAMA results have been corrected using the relationship shown in Figure 16. Additionally, DESI BGS results after correcting for the Local Void bias are shown (purple dots). This correction is based on 2019ApJ...872..180C, leveraging the ELUCID simulation (2014ApJ...794...94W; 2016ApJ...831..164W), which uses the reconstructed initial density field of the nearby universe. The GSMF from (2022MNRAS.513..439D) for Local Group galaxies (brown triages; 2012AJ....144....4M) is also reproduced using the same stellar mass calibration as applied to GAMA.
@@
Figure 16: Left: Comparing the stellar mass $M_{*}^{\rm{GAMA}}$ obtained from GAMA DR4 (2022MNRAS.513..439D) with those derived using our CIGALE SED template (described in Section 2.1) based on DECaLS grzgrz bands photometry ($M_{*}^{\rm{CIGALE}}$). The relationship between the two can be effectively characterized by a power law: $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}{\rm M}_{\sun})=1.072\log_{10}(M^{\rm{GAMA}}_{*}/10^{10}{\rm M}_{\sun})-0.237$ (red line). Right: Comparing the stellar mass $M_{*}^{\rm{COSMOS}}$ obtained from COSMOS2020 (2023A&A...677A.184W) with those derived using our CIGALE SED template based on HSC grzgrz bands photometry at $0.2<z_{\rm p}<0.8$. Black dots with errorbars represent the mean relation with the 1 $\sigma$ scatter. The relationship between the two can be effectively characterized by a power law: $\log_{10}(M^{\rm{CIGALE}}_{*}/10^{10}{\rm M}_{\sun})=0.948\log_{10}(M^{\rm{COSMOS}}_{*}/10^{10}{\rm M}_{\sun})-0.012$.
@@
Figure 17: Top: Comparison with galaxy stellar mass functions (GSMFs) from the COSMOS field at higher redshifts. The GSMF of the full sample (black dots) is compared with results from COSMOS-Web (2025A&A...695A..20S) at $0.2<z_{\rm p}<0.5$ (blue dots) and $0.5<z_{\rm p}<0.8$ (pink dots). Stellar masses from COSMOS-Web have been calibrated to match our scale using the right panel of Figure 16. GSMFs from G10-COSMOS (2017MNRAS.470..283W) at $z_{\rm p}<1$ are also shown (green triangles) with stellar mass calibrated with the relation for GAMA. Bottom: Stellar mass density within each stellar mass interval.